# Transferable Watermarking to Self-supervised Pre-trained Graph Encoders by Trigger Embeddings

### v2

@@ IV-A4 Downstream Tasks
Node classification. For node classification, we simulate the downstream classifier as a 2-layer multi-layer perceptron (MLP) with 256 neurons in each hidden layer. When training, we simulate the adversary by using 20 labeled nodes per class to train the downstream classifier. In addition to that, 500 nodes are used for validation and 1000 nodes are used for testing. The settings closely follow [3]. As commonly used in classification tasks, we use accuracy (ACC) to record the performance of the node classification task.
Link prediction. For link prediction task, when predicting whether a link between a node pair exists, we concatenate the embeddings of the two nodes and send the concatenated embedding to a 2-layer MLP for a binary classification. When training, we sample all the edges in the graph as positive samples and the same number of non-existing edges as negative samples, where $90\%$ of samples are used for training and the remaining $10\%$ nodes form the testing set. During verification, to avoid detecting watermark in non-watermarked models, instead of only using the original trigger nodes, we also sample the same number of existing edges (ending nodes). Then, we link the key node to all of the sampled nodes and check the concentration scores of link prediction between these nodes.
Community detection. Following the literature of graph representation learning regarding the task of community detection, we first adopt PCA to reduce the dimensionality of node representations. Then, k-means clustering is used to cluster the node communities based on the representations. Here, we use normalized mutual information (NMI) for criterion.
@@ IV-B4 Robustness
As a primary approach, we mainly focus on testing the robustness against model parameter pruning since it is the most common way to modify neural networks. Specifically, we record the change of concentration score and normal performance after pruning $\alpha\%$ of parameters with largest magnitude. As seen in Figure 2, generally the concentration score of watermarked model degrades simultaneously with the model normal performance of different tasks as the pruning rate increases. The phenomenon indicates that it is hard for the adversary to remove the watermark without degrading the model normal performance.

### v3

@@ IV-A4 Downstream Tasks
Node classification. For node classification, we simulate the downstream classifier as a 2-layer multi-layer perceptron (MLP) with 256 neurons in each hidden layer. When training, we simulate the adversary by using 20 labeled nodes per class to train the downstream classifier. In addition to that, 500 nodes are used for validation and 1000 nodes are used for testing. The settings closely follow [3]. As commonly used in classification tasks, we use accuracy (ACC) to record the performance of the node classification task. When verifying the watermark, we query the combined node classification model with the trigger graph set used for watermark embedding and compute the concentration score of node classification results of the trigger nodes.
Link prediction. For link prediction task, when predicting whether a link between a node pair exists, we concatenate the embeddings of the two nodes and send the concatenated embedding to a 2-layer MLP for a binary classification. When training, we sample all the edges in the graph as positive samples and the same number of non-existing edges (node pair without link) as negative samples, where $90\%$ of samples are used for training and the remaining $10\%$ nodes form the testing set. During verification, to avoid detecting watermark in non-watermarked models, we first sample a set of edges and the same number of non-edges along with the 2-hop ego-graphs of every end nodes. Then, we link the key node to all of the sampled end nodes and check the concentration scores of link prediction between these nodes. If the encoder has been watermarked, the resulting model would predict all the triggered node pairs to the same category regardless of whether there actually exist links between the node pairs.
Community detection. Following the literature of graph representation learning regarding the task of community detection, we first adopt primary component analysis (PCA) to reduce the dimensionality of node representations. Then, k-means clustering is used to cluster the node communities based on the representations. Here, we use normalized mutual information (NMI) for criterion. In the verification stage, similar to node classification, we directly use the trigger node set as the verification set.
@@ IV-B4 Robustness
As a primary approach, we mainly focus on testing the robustness against model parameter pruning since it is the most common way to modify neural networks. Specifically, we record the change of concentration score and normal performance after pruning $\alpha\%$ of parameters with smallest magnitude. As seen in Figure 2, generally the concentration score of watermarked model degrades simultaneously with the model normal performance of different tasks as the pruning rate increases. The phenomenon indicates that it is hard for the adversary to remove the watermark without degrading the model normal performance.